# Pose2Gaze: Generating Realistic Human Gaze Behaviour from Full-body Poses using an Eye-body Coordination Model

### v1

@@ 5.2 Gaze Generation Results
Table 4 summarises the mean angular errors of different methods for generating eye gaze from future body poses. We can see that for pick, place and all actions on the MoGaze dataset, our method consistently outperforms the state of the art, achieving an average improvement of $25.8\%$ ($9.8^{\circ}$ vs. $13.2^{\circ}$), $26.5\%$ ($8.6^{\circ}$ vs. $11.7^{\circ}$), and $26.4\%$ ($9.2^{\circ}$ vs. $12.5^{\circ}$) respectively. On the GIMO dataset, our method outperforms the state of the art with an average improvement of $21.6\%$ ($15.2^{\circ}$ vs. $19.4^{\circ}$) in terms of mean angular error. A paired Wilcoxon signed-rank test was conducted and the results validated that the differences between our method and the state of the art are statistically significant ($p<0.01$). The above results demonstrate that our method has high performance in generating eye gaze from future body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{9.8}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
MoGaze-place $\textbf{8.6}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
MoGaze-all $\textbf{9.2}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
GIMO-all $\textbf{15.2}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
Table 4: Mean angular errors of different methods for generating eye gaze from future poses on the MoGaze and GIMO datasets.
5.3 Summary
The results in above section demonstrate that our method significantly outperforms the state-of-the-art methods for three different generation tasks. Furthermore, by comparing our method’s performances under different generation settings, we find that our method achieves its best performance in using future body poses ($9.2^{\circ}$ on MoGaze and $15.2^{\circ}$ on GIMO), followed by using present body poses ($9.9^{\circ}$ on MoGaze and $16.0^{\circ}$ on GIMO) and then using past poses ($13.1^{\circ}$ on MoGaze and $18.6^{\circ}$ on GIMO). This result corresponds with our analysis in subsection 3.3 that eye gaze has highest correlations with body motions in the near future and suggests that in offline applications where future pose is available, e.g. creating and animating virtual humans [3, 41], using future body poses could generate more realistic eye gaze.
@@ 5.3 Summary
5.4 Ablation Study
We finally performed an ablation study to evaluate the effectiveness of each component used in our model.
Effectiveness of Our Model Architecture
@@ 5.4 Ablation Study
To evaluate the effectiveness of our model architecture, we evaluated different ablated versions of it that did not contain the DCT, spatial GCN, temporal GCN, the body poses input, and the head directions input, and re-trained the ablated models. Table 5 summarises the performances of different ablated versions of our method for generating eye gaze on the GIMO dataset. We can see that our method consistently outperforms the ablated methods under different generation settings and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$), thus underlining the effectiveness of our model architecture. One exception is that our method is not as good as the w/o DCT and w/o T-GCN versions in generating eye gaze from past body poses. This is probably because past body poses have weaker correlations with eye gaze than present and future body poses (subsection 3.3), and thus the temporal features extracted from past body poses using DCT and temporal GCN are not as effective as that extracted from present and future poses.
Ours w/o DCT w/o S-GCN w/o T-GCN w/o Poses w/o Head
Past $18.6^{\circ}$ $\textbf{18.2}^{\circ}$ $19.0^{\circ}$ $18.3^{\circ}$ $22.8^{\circ}$ $23.2^{\circ}$
Present $\textbf{16.0}^{\circ}$ $16.6^{\circ}$ $17.3^{\circ}$ $16.1^{\circ}$ $20.7^{\circ}$ $20.4^{\circ}$
Future $\textbf{15.2}^{\circ}$ $16.4^{\circ}$ $17.2^{\circ}$ $16.0^{\circ}$ $20.7^{\circ}$ $19.5^{\circ}$
Table 5: Mean angular errors of different ablated versions of our method for generating eye gaze under different settings on GIMO.
Importance of Different Body Parts
We further evaluated the importance of different body parts for eye gaze generation. To this end, we re-trained our model using different body parts as input, i.e. only using the Torso part, Arm part, Leg part, as well as different combinations of these. Table 6 shows the performances of our method for generating eye gaze from different body parts on the GIMO dataset. As can be seen from the table, our method achieves significantly better performances than other versions (paired Wilcoxon signed-rank test, $p<0.01$), validating that each body part used in our model can help improve the performance of eye gaze generation.
Ours Torso Arm Leg Torso+Arm Torso+Leg Arm+Leg
Past $\textbf{18.6}^{\circ}$ $20.7^{\circ}$ $19.4^{\circ}$ $19.7^{\circ}$ $19.6^{\circ}$ $19.6^{\circ}$ $18.8^{\circ}$
Present $\textbf{16.0}^{\circ}$ $17.8^{\circ}$ $17.4^{\circ}$ $18.0^{\circ}$ $16.2^{\circ}$ $17.0^{\circ}$ $16.9^{\circ}$
Future $\textbf{15.2}^{\circ}$ $18.6^{\circ}$ $16.4^{\circ}$ $18.3^{\circ}$ $16.0^{\circ}$ $16.7^{\circ}$ $16.7^{\circ}$
Table 6: Mean angular errors of our method for generating eye gaze from different body parts under different settings on GIMO.
6 Discussion
@@
Our analyses on eye-body coordination in daily activities effectively guided the design of our model. Specifically, in the analysis of correlation between eye gaze and body orientations, we find that eye gaze has strong correlations with the orientations of different body joints, especially with head directions, and these body orientations are strongly correlated with each other (Table 1). Inspired by this, our model extracts body orientation features directly from head directions rather than all the body orientations (subsection 4.1) because head direction itself contains sufficient information for generating eye gaze. We also find that there exists little or no time delay between head and eye movements (Figure 1). Therefore, in the task of generating gaze from future poses (subsection 5.1), we used head directions at the present time to generate target eye gaze rather than using future head directions. In the analysis of correlation between eye gaze and body motions, we find that eye gaze is strongly correlated with full-body motions (Figure 2). Therefore, our model extracts features from human full-body poses to generate realistic eye gaze (subsection 4.2). We also find that there exists a noticeable time delay between full-body motions and eye gaze (Figure 3). Inspired by this, we used future body poses as input to improve the performance of eye gaze generation in situations where future pose information is available (subsection 5.2). The good performance of our model validated that our analyses are effective and significant for the design of eye-body coordination models.
Pose-based Gaze Generation
Prior methods have typically generated eye gaze from head movements while our method generates eye gaze from full-body poses. Our method significantly outperforms prior methods on both the MoGaze and GIMO datasets for three different generation tasks (Table 2, Table 3, Table 4) and the ablation studies validated that each body part contributes to the performance of our method (Table 6). These results reveal the significant potential of human full-body poses for generating realistic eye movements and thus open the promising research direction of pose-based gaze generation.
@@
We evaluated our method for three different generation tasks that are respectively linked to different kinds of applications (subsection 5.1). The fact that our method outperforms the state-of-the-art methods by a large margin under all the three different generation settings (Table 2, Table 3, Table 4) demonstrates that our method has great potential for these numerous applications including dynamic event triggering [14], gaze-contingent rendering [21, 20], gaze-based interaction [9], as well as creating and animating virtual humans [3, 41]. In addition, with the development of motion capture and motion generation techniques [3, 15], it becomes increasingly easy to obtain human full-body poses. Our approach could be integrated with such motion capture and motion generation techniques to generate realistic eye and full-body movements simultaneously.
Limitations
Despite these advances, we identified several limitations of our work that we plan to address in the future. First, we conducted our analysis and experiments using two public indoor datasets that contain eye and full-body movements in various daily activities. However, there exist some other interesting activities that are not covered by these datasets, e.g., narration or conversation actions [3]. No suitable gaze dataset exists yet, so the coordination of eye and full-body movements for these activities still remains to be explored. In addition, all activities in our datasets were performed in indoor environments and it remains to be seen whether eye-body coordination as characterised here also applies to outdoor scenarios. Furthermore, the activities we studied only cover the interaction between humans and objects and don’t involve other interaction paradigms, e.g. interaction between two humans and interaction between a human and a social robot. Eye-body coordination in such interaction paradigms may be different and deserves to be studied in future work.
Future Work
Besides overcoming the above limitations, many potential avenues of future work exist. First, it is well-known that human eye gaze is influenced by both the bottom-up scene content and the top-down tasks. Therefore, it will be interesting to analyse how the bottom-up and top-down factors influence the eye-body coordination. In addition, incorporating other modalities such as audio, text, or speech into our model may further boost our model’s gaze generation performance. Finally, generating stylistic eye gaze, e.g. eye gaze that can convey different emotions [39], is an interesting avenue to extend our work.
7 Conclusion
@@
In this work we were the first to explore the challenging task of generating realistic human eye gaze from human full-body poses. We first conducted a comprehensive analysis on the coordination of human eye and full-body movements in everyday activities and revealed that eye gaze is strongly correlated with head directions and also full-body motions. Based on these insights, we proposed a novel eye-body coordination model to generate eye gaze from head directions and full-body poses that outperformed the state-of-the-art methods by a large margin for three different generation tasks. Taken together, our work provides novel insights into eye-body coordination during daily activities and makes an important step towards more holistic, cross-modal generation of virtual agent behaviour.

### v3

@@ V-B Gaze Generation Results
Table IV-future summarises the mean angular errors of different methods for generating eye gaze from future body poses on the MoGaze, ADT, GIMO, and EgoBody datasets. We can see that our method outperforms prior methods in both individual actions and all the actions in a dataset, achieving an average improvement of $24.0\%$ ($9.5^{\circ}$ vs. $12.5^{\circ}$) on MoGaze, $10.1\%$ ($8.9^{\circ}$ vs. $9.9^{\circ}$) on ADT, $21.3\%$ ($15.5^{\circ}$ vs. $19.7^{\circ}$) on GIMO, and $25.4\%$ ($12.9^{\circ}$ vs. $17.3^{\circ}$) on EgoBody. A paired Wilcoxon signed-rank test was conducted and the results validated that the differences between our method and the state of the art are statistically significant ($p<0.01$) in all the four datasets. The above results demonstrate that our method has high performance in generating eye gaze from future body poses in both human-object and human-human interaction activities.
Summary
The above results demonstrate that our method significantly outperforms the state-of-the-art methods for three different eye gaze generation tasks. Furthermore, we find that in human-object interaction activities, our method achieves significantly higher accuracies using future body poses than using past body poses: $9.5^{\circ}$ vs. $13.1^{\circ}$ on MoGaze, $8.9^{\circ}$ vs. $11.5^{\circ}$ on ADT, and $15.5^{\circ}$ vs. $18.4^{\circ}$ on GIMO. These results correspond with our analysis in subsection III-C that eye gaze has highest correlations with body motions in the near future and suggest that in offline applications where future pose is available, e.g. virtual content design and optimisation [9, 5], using future body poses could generate more accurate eye gaze directions. In human-human interaction activities, we find that our method using past, present, or future body poses achieve similar performances: $13.2^{\circ}$ vs. $13.0^{\circ}$ vs. $12.9^{\circ}$ on EgoBody. This is because the eye-body correlations in human-human interactions are very close at $-500$ $ms$ (past), 00 $ms$ (present) or $500$ $ms$ (future), as illustrated in Figure 3b.
Model Size and Time Costs
@@
V-C Ablation Study
@@
We finally performed an ablation study to evaluate the effectiveness of each component used in our model. Specifically, we respectively evaluated different ablated versions of our model that did not contain the DCT, spatial GCN, temporal GCN, the body poses of the person, the body poses of the interaction partner, and the head directions under three generation settings. Table V summarises the mean angular errors of different ablated versions of our method on the ADT, GIMO, and EgoBody datasets. We can see that our method consistently outperforms the ablated methods under different generation settings and the results are statistically significant (paired Wilcoxon signed-rank test, $p<0.01$), thus underlining the effectiveness of our model architecture. In addition, we find that in human-human interactions removing the body poses of the person or the body poses of the interaction partner significantly degrades the performance of our model: EgoBody-past from $13.2^{\circ}$ to $20.6^{\circ}$ or $18.6^{\circ}$, EgoBody-present from $13.0^{\circ}$ to $18.1^{\circ}$ or $17.9^{\circ}$, EgoBody-future from $12.9^{\circ}$ to $17.1^{\circ}$ or $17.9^{\circ}$. These results demonstrate the effectiveness of the two body poses on generating eye gaze in human-human interaction activities.
TABLE V: Mean angular errors of different ablated versions of our method on the ADT, GIMO, and EgoBody datasets. Pose refers to the body pose of the person while Pose_I means the body pose of the interaction partner.
Ours w/o DCT w/o S-GCN w/o T-GCN w/o Pose w/o Pose_I w/o Head
ADT past $\textbf{11.5}^{\circ}$ $11.7^{\circ}$ $11.8^{\circ}$ $11.9^{\circ}$ $12.2^{\circ}$ $18.2^{\circ}$
present $\textbf{9.0}^{\circ}$ $9.1^{\circ}$ $9.4^{\circ}$ $9.1^{\circ}$ $9.5^{\circ}$ $17.7^{\circ}$
future $\textbf{8.9}^{\circ}$ $9.1^{\circ}$ $9.3^{\circ}$ $9.1^{\circ}$ $9.3^{\circ}$ $16.4^{\circ}$
GIMO past $\textbf{18.4}^{\circ}$ $19.0^{\circ}$ $19.3^{\circ}$ $19.1^{\circ}$ $21.2^{\circ}$ $22.1^{\circ}$
present $\textbf{16.3}^{\circ}$ $17.3^{\circ}$ $18.1^{\circ}$ $17.3^{\circ}$ $20.8^{\circ}$ $20.9^{\circ}$
future $\textbf{15.5}^{\circ}$ $16.6^{\circ}$ $18.1^{\circ}$ $16.7^{\circ}$ $20.8^{\circ}$ $18.8^{\circ}$
EgoBody past $\textbf{13.2}^{\circ}$ $13.5^{\circ}$ $13.4^{\circ}$ $13.4^{\circ}$ $20.6^{\circ}$ $18.6^{\circ}$ $15.1^{\circ}$
present $\textbf{13.0}^{\circ}$ $13.1^{\circ}$ $14.3^{\circ}$ $13.3^{\circ}$ $18.1^{\circ}$ $17.9^{\circ}$ $14.5^{\circ}$
future $\textbf{12.9}^{\circ}$ $13.7^{\circ}$ $14.8^{\circ}$ $13.5^{\circ}$ $17.1^{\circ}$ $17.9^{\circ}$ $15.1^{\circ}$
VI Eye-based Activity Recognition
Activity recognition is a popular task in the area of human-centred computing and has great relevance for many VR/AR scenarios including adaptive virtual environment design [2], low-latency predictive interfaces [45, 46], or human-aware intelligent systems [47]. It is well-known that user activities can be recognised directly from human eye gaze [7, 48]. As such, eye-based activity recognition serves as a particularly relevant sample downstream task to further evaluate our gaze prediction method. The higher the activity recognition accuracy, the better the quality of the predicted gaze sequences.
Datasets
We only evaluated on the ADT and EgoBody datasets given that they provide activity labels for each recorded sequence and their sequence lengths are sufficient ($>$ 10 s) for training activity recognition methods [7, 49]. For both datasets, we used the same training and test sets as in section V.
Activity Recognition Method
We used the state-of-the-art method EHTask [7] for activity recognition in VR. EHTask takes a sequence of eye gaze data as input, extracts eye features using a 1D CNN and a bidirectional gated recurrent unit (GRU), and finally recognises user activities from the eye features using fully-connected layers.
Procedure
We first trained EHTask to recognise user activities from the ground truth eye gaze sequences using default parameters. On the ADT dataset, we trained EHTask to recognise three activities: decoration, meal, and work. On the EgoBody dataset, we trained EHTask for three activities that have the most recordings, i.e. chat, learn, and teach. Since activity recognition usually requires a long sequence as input [7, 49], we used 300 frames (corresponding to 10 seconds) of ground truth eye gaze data to train EHTask as in the original paper [7]. At test time, we used the eye gaze sequences generated from different methods as input to EHTask and evaluated its activity recognition performance. Since the gaze prediction methods were trained to predict 15 frames of eye gaze while EHTask requires an input of 300 frames, we first segmented the 300 frames of test data into 20 windows (each window contains 15 frames), then predicted eye gaze for each window, and finally concatenated the gaze predictions of all the windows and used them as input to EHTask to recognise user activities.
Results
Table VI shows the recognition performances of using the ground truth eye gaze and the eye gaze predicted from different methods on ADT and EgoBody. It can be seen in the table that our method achieves higher recognition accuracies than other methods on both ADT ($70.0\%$ vs. $67.3\%$) and EgoBody ($60.1\%$ vs. $58.2\%$). Using a paired Wilcoxon signed-rank test we confirmed that these differences were statistically significant ($p<0.01$). We also notice that our method can achieve comparable performance with the ground truth eye gaze in terms of recognition accuracy ($70.0\%$ vs. $74.7\%$ on ADT, and $60.1\%$ vs. $62.1\%$ on EgoBody). These results illustrate the benefit of our method when using the predicted gaze sequences for downstream tasks with high relevance for VR/AR applications, such as eye-based activity recognition.
TABLE VI: Eye-based activity recognition accuracies of using ground truth eye gaze and the eye gaze predicted from different methods on ADT and EgoBody. Best results are in bold while the second best are underlined.
GT Ours DGaze[11] FixationNet[1] Head Direction Chance
ADT 74.7% 70.0% 67.3% 66.8% 40.9% 33.3%
EgoBody 62.1% 60.1% 52.3% 58.2% 50.3% 33.3%
VII Discussion
@@
Our analyses on eye-body coordination in daily activities revealed novel insights for understanding human eye and body movements that we also used to guide the design of our eye-body coordination model. Specifically, in the analysis of correlation between gaze and head directions, we found that there exists little or no time delay between head and eye movements (see Figure 1). Therefore, in the task of generating gaze from future poses (see subsection V-A), we used head directions at the present time to generate target eye gaze rather than using future head directions. In the analysis of correlation between eye gaze and body motions, we found that gaze is strongly correlated with full-body motions in human-object interactions (see subsection III-C) and has high correlations with the directions between two bodies in human-human interactions (see subsection III-D). Therefore, our model uses features from the body pose of the person to generate gaze in human-object interactions and employed the features from the body poses of the person and the interaction partner to predict gaze in human-human interactions (see Figure 4). We also found that eye movements precede body motions in human-object interactions (see subsection III-C). Exploiting this, we used future body poses as input to improve the performance of eye gaze generation in situations where future pose information is available (see subsection V-B). The superior performance of our model shows that our analyses are effective and are valuable for the design of future eye-body coordination models.
Pose-based Gaze Prediction
Prior methods have typically generated eye gaze from head movements while we predict eye gaze from both head movements and full-body poses. This approach proved highly effective and significantly outperforms prior methods in both human-object and human-human interaction activities for three different eye gaze generation tasks (see Table IV). An ablation study confirmed that body poses are instrumental for achieving these performance improvements (see Table V). Taken together, these results underline the significant potential of human full-body poses for generating eye movements and thus open the promising research direction of pose-based gaze prediction.
@@
In addition to the sample downstream task of eye-based activity recognition (see section VI), we believe that our method has significant potential for numerous VR/AR applications, including dynamic event triggering [2], gaze-contingent rendering [10, 11], gaze-based interaction [24], or virtual content design and optimisation [9, 5]. In addition, with further advances in the development of motion capture and motion generation techniques [19, 16], it will become increasingly easy to obtain human full-body poses in VR and AR environments. In the future, our approach could be integrated with such motion capture and motion generation techniques to generate eye and full-body movements simultaneously.
Limitations and Future Work
Despite these advances, our work has several limitations that we plan to address in the future. First, we focused on eye-body coordination in different activities and ignored the influence of the visual scene in front of the person. It will be interesting to explore whether eye-body coordination during the same activity will be affected by different scene content. If so, this will also open a new avenue for research on ways to incorporate visual scene information into the prediction and to see whether this can improve performance further. Second, incorporating other modalities such as facial expressions and audio signals in social interactions may also boost our model’s gaze prediction performance. In addition, existing datasets that include body pose and eye gaze have all been collected in indoor environments, thus unfortunately limiting our analyses to indoor settings. It remains to be seen whether eye-body coordination as characterised here also generalises to outdoor scenarios. Furthermore, existing datasets only offer data for interactions between two humans, thus limiting our analyses to dyadic settings. It will be interesting to explore the eye-body coordination during other interaction settings that are important for VR/AR applications, e.g. interactions between more humans and interactions between a human and a virtual avatar. Finally, generating stylistic eye gaze, e.g. eye gaze that can convey different emotions [32], is an interesting avenue for extending our work in the future.
VIII Conclusion
@@
In this work we were the first to explore the challenging task of predicting human eye gaze from full-body poses. We first conducted a comprehensive analysis on the coordination of human eye and full-body movements in everyday activities and revealed that in human-object interactions eye gaze is strongly correlated with full-body motions while in human-human interactions a person’s gaze direction has strong correlations with the directions pointing from his body to the body of the interaction partner. Based on these insights, we proposed a novel eye-body coordination model to predict eye gaze from head directions and full-body poses that outperformed the state-of-the-art methods by a large margin for three different prediction tasks as well as the application of eye-based activity recognition. Taken together, our work provides novel insights into eye-body coordination during daily activities and makes an important step towards pose-based gaze prediction.
Acknowledgement
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2075 – 390740016. A. Bulling was funded by the European Research Council (ERC) under grant agreement 801708.